OpenStudy (anonymous):

What is the mode of this set of data? 1982 1988 1989 1994 1995 2005 Answer 1988 1989 11994 This data set has no mode. Question 7 What is the median of this set of data: 10, 11, 12, 12, 15, 19, 20, 21, and 22? Express your answer as an integer. Answer 12 15 16 19 Question 8 What is the mean of this set of data: 78, 79, 80, 80, 89, 89, 92, 93, 94, 95, 95, 95? Express your answer as a decimal number rounded to the nearest tenth. Answer 88.3 89 92 88.6

OpenStudy (shinalcantara):

Mode is the data with the most number of appearances. Or the most frequent value. Say you have the set of data as 1,5,7,8,8,9 In this case, 8 is the mode since it appeared twice

OpenStudy (anonymous):

omg thank you so much i actually understand it now but can you please help me with the other questions thnx so much

OpenStudy (shinalcantara):

Median is the middle of the given set Taking the same set: 1,5,7,8,8,9 Notice that there are two middle values-7and8 then take its average. (7+8)/2 = 7.5 Therefore the median is 7.5

OpenStudy (shinalcantara):

For this set of data 1,5,7,8,9 The middle term is 7. Therefore 7 is the median

OpenStudy (anonymous):

is the answer to Q2 15?

OpenStudy (shinalcantara):

Mean simply implies that you need to get the average value of the set of data. Say you have the same set of data as above 1,5,7,8,8,9 Get the sum of it 1+5+7+8+8+9 = 38 Divide it by the number of data being added. In this case there are 6 of them so you'll have 38/6 = 6.33 6.33 is the mean

OpenStudy (shinalcantara):

yep it's 15

OpenStudy (anonymous):

thank you so much i actually understand now *high five* lol

OpenStudy (shinalcantara):

yw :)

OpenStudy (anonymous):

change ur name to (smart) lol its a perfect name for you

Directrix (directrix):

@Awesomemeow What did you get for the mean of 78, 79, 80, 80, 89, 89, 92, 93, 94, 95, 95,95 ?